OpenStudy (anonymous):

how to convert 13.1 MeV to kj/mol?

OpenStudy (frostbite):

Lets write up a few things: \[1 eV=1,60217653 * 10^{-19}J\] \[MeV=1000000 eV =10^{6} eV\] \[1 kJ=1000 J\] \[N _{a}=6,02214*10^{23} \frac{ 1 }{ mol }\] So: \[13.1MeV*10^{6} \frac{ eV }{ MeV }*1,60217653 * 10^{-19} \frac{ J }{ eV }*10^{-3} \frac{ kJ }{ J }*6,02214*10^{23} \frac{ 1 }{ mol }\]